If P = $8 and MC = $5 + 2Q, the competitive firm's profit-maximizing level of output is:

a.

1.5

b.

0.2

c.

8

d.

15

     2.   If a firm faces the demand curve P = 60 – Q and the price is $30, the consumer surplus is:

a.

200

b.

300

c.

450

d.

650

     3.   If the demand curve is P = 60 – Q and the supply curve is Q = P, the market equilibrium output is:

a.

45

b.

30

c.

60

d.

15

     4.   If the demand curve is P = 60 – Q and the supply curve is Q = P, the total revenue of the firm at equilibrium is:

a.

300

b.

500

c.

750

d.

900
Answer #1


Question 1

A competitive firm maximizes profit when it produce that level of output corresponding to which price equals the marginal cost.

Equating the price and the marginal cost -

P = MC

8 = 5 + 2Q

2Q = 8 - 5

2Q = 3

Q = 3/2 = 1.5

The competitive firm's profit-maximizing level of output is 1.5 units.

Hence, the correct answer is the option (a).

Question 2

Demand curve is as follows -

P = 60 - Q

Calculate the price when demand is zero -

P = 60 - Q = 60 - 0 = 60

The given market price is $30.

Calculate the demand at the market price -

P = 60 - Q

30 = 60 - Q

Q = 30

Calculate the consumer surplus -

CS = 1/2 * (Price when demand is zero - Market price) * Quantity demanded at the market price

CS = 1/2 * (60 - 30) * 30

CS = 1/2 * 30 * 30 = 450

The consumer surplus is $450.

Hence, the correct answer is option (c).

Question 3

Demand curve is as follows -

P = 60 - Q

Supply curve is as follows -

Q = P

or,

P = Q

At the market equilibrium, demand equals supply

So,

60 - Q = Q

2Q = 60

Q = 60/2 = 30

Thus,

The market equilibrium output is 30 units.

Hence, the correct answer is the option (b).

Question 4

Demand curve is as follows -

P = 60 - Q

Supply curve is as follows -

Q = P

or,

P = Q

At the market equilibrium, demand equals supply

So,

60 - Q = Q

2Q = 60

Q = 60/2 = 30

Thus,

The market equilibrium output is 30 units.

P = 60 - Q = 60 - 30 = 30

The market equilibrium price is $30 per unit.

Calculate the Total revenue -

Total revenue = Price * Quantity = 30 units * $30 = $900

Thus,

The Total revenue is $900.

Hence, the correct answer is the option (d).